OpenStudy (anonymous):

How do I solve 4 pi=w-6r?

OpenStudy (anonymous):

What are you attempting to solve for? If w, \[w = 4\pi + 6r \] If r, \[6r + 4\pi = w\\6r = w - 4\pi\\r = \frac{w - 4\pi}{6} \]